OpenStudy (anonymous):

Determine where f(x) intersects the x-axis. f(x)=x2−18x+80

OpenStudy (anonymous):

f(x) intercepts the x-axis when y=0. x² - 18x + 80 = 0 if you find the roots, you will find your answer.

OpenStudy (anonymous):

(x+10)(x-8)?

OpenStudy (anonymous):

Notice that -10 -8 = - 18 and that (-10)(-8) = +80. So, the factoring would be (x-10)(x-8) = 0, what gives us x = 10 and x = 8.

OpenStudy (anonymous):

Ahh okay my mistake, so factoring it out would be (x-10)(x-8) was the last part of your response a question? What is the next step from here? thank you for your help

OpenStudy (anonymous):

If x = 10 and x = 8, and you already know that you made y = 0, what are the coordinates? First one is (10,0) and the second? Anytime.